OpenStudy (anonymous):

Identify the solutions of 3x^2-13x+4=0 A. x=-1/3, x=-4 B. x=1/3, x=4 C. x=-4/3, x=-1 D. x=4/3, x=1

OpenStudy (jack1):

y= ax^2 + bx +c in your case a =3 b= -13 c = 4

OpenStudy (jack1):

do you know how to factor this equation?

OpenStudy (anonymous):

I dont know how to do any of it ._. It's all friggin confusing as hell to me...

OpenStudy (jack1):

cool so we'll start there then: factors of "a" what times what = 3?

OpenStudy (anonymous):

1 times 3? lol

OpenStudy (jack1):

perfect! so we'll have something that looks like : ( 3x + __ ) (1x + __ )

OpenStudy (jack1):

now, the factors of "c" what times what = 4 (need both answers here)

OpenStudy (anonymous):

1 times 4 2 times 2

OpenStudy (jack1):

easy as, we're breezin through

OpenStudy (jack1):

now to pick which one is ours, it has to add to equal "b" (-13) also, as its a negative, both our factors are probably going to be negatives so as "a" was 3 and 1: 3 times (4 or 1 or 2) + (4 or 1 or 2) = - 13

OpenStudy (jack1):

so pick your poison: -1,-4 or -2,-2 personally i'd go with the 3 times -4 (this equals -12) and -1 as this matches perfectly

OpenStudy (jack1):

so we've picked the factors that match this equation so y = ax^2 + bx +c y = 3x^2 -13x +4 y = (3x -1) (x - 4) and as y = 0...

OpenStudy (jack1):

0 = (3x -1) (x - 4) therefore what times what = 0? answer: anything times 0 =0 so either 3x - 1 = 0 or x - 4 = 0 so either 3x = 1 or x = 4 so either x = 1/3 or x = 4

OpenStudy (jack1):

so one of those answers above matches up with those values of x

OpenStudy (anonymous):

im lost ._. I still don't get it...

OpenStudy (anonymous):

FML ill just guess. >.<

OpenStudy (jack1):

so either x = 1/3 or x = 4

OpenStudy (jack1):

both of these answers are correct, the graph of the line crosses the x axis at x=1/3 and again at x = 4, so your answer is b